A man is 42 years old and his son is 12. In how many years will the father be twice as old as his son.
Answer by unlockmath(1599)

Hello,
We can set up an equation where x represents the years.
42+x=(12+x)2
Expand:
42+x=24+2x
Subtract x and 24:
x=18 years
Make sense?
RJ
www.math-unlock.com